# Learning Patterns from Biological Networks: A Compounded Burr Probability Model

### v2

@@ III Model Formulation
As a practical example, we consider a complex biological network, such as a protein-protein interaction network, which consists of various interacting proteins. In this example, the network undergoes dynamic processes, and we want to investigate the time it takes for a specific protein to be activated or inhibited, denoted as the activation/inhibition time. Each vertex in the network represents a specific entity, such as a protein in a protein-protein interaction network. The activation/inhibition time for a specific protein (vertex) is defined as the minimum value among the random variables $X_{i}$, which represents the time intervals between these events. Assuming that the activation/inhibition events are stochastic occurrences that can happen multiple times in the network’s dynamics, we model the time interval between each event using independent and identically distributed random variables $X_{i}$, where $i=1,2,\cdots,N$ and $(N-1)$ is a Poisson random variable with parameter $0\leq\lambda\leq\infty\;$ Erdos, Rényi et al. (1960); Marshall and Olkin (1997). The continuous probability distribution function for these random variables is represented as $F(\cdot)$. Therefore, the activation/inhibition time for the specific protein of interest is defined as the minimum value among these random variables:
$Y=\min\{X_{1},X_{2},\cdots,X_{N}\}.$
Here, $Y$ represents the activation/inhibition time for the specific protein of interest, and $N$ is the number of such events (activation/inhibition events). Let us assume that $N-1$ is independent of the random variables $X_{i}$ and follows a Poisson distribution with parameter $\lambda$, where $0\leq\lambda<\infty$. In this scenario, we can determine the probability mass function of the number of activation or inhibition events, denoted as $N-1$, which is given by:
$P(N-1=n-1)=\frac{e^{-\lambda}\lambda^{n-1}}{(n-1)!},\;n=1,2,...$
where the parameter $\lambda$ determines the expected number of these activation or inhibition events occurrences over a specified period. Further, the survival function of the minimum activation/inhibition time, denoted as $Y$, can be calculated as
$\bar{G}(y)$ $=P\left[X_{1}\geq y,X_{2}\geq y,\ldots,X_{n}\geq y\right]:=P(Y\geqslant y)$ (1)
@@
$=\displaystyle\sum_{n=1}^{\infty}\left[\bar{F}(y)\right]^{n}\frac{e^{-\lambda}\lambda^{n-1}}{(n-1)!}=\frac{\bar{F}(y)}{\exp\{\lambda F(y)\}},$
where $\bar{F}(y)$ represents the survival function of $X$. It is worth noting that the ratio ${\bar{F}(y)}/{\bar{G}(y)}$ is non-decreasing for $y\geqslant 0$, shown as
@@
Using the same arguments as discussed above, we can calculate the maximum activation/inhibition time, which can be defined as
$Z=\max\{X_{1},X_{2},\cdots,X_{N}\}.$
The distribution function of the maximum activation/inhibition time, denoted as $Z$, can be obtained as
@@
This method acts as a skewing mechanism, as illustrated by the following theorem:
@@ III.1 Hazard rates and risk functions
In the context of reliability and survival data analysis, we introduce a transformation to the distribution $G(y:\lambda)$, which incorporates a new parameter $\lambda\geq 0$ into any existing family of continuous distributions. This transformation is defined based on the survival function of the distribution given:
$\bar{G}(y)$ $=\frac{\bar{F}(y)}{\exp\{\lambda F(y)\}}.$
By assuming the continuity of $F$ throughout, we can derive the corresponding density function as follows:
$g(y;\lambda)$ $=\frac{f(y)\left[1+\lambda\bar{F}(y)\right]}{\exp\{\lambda F(y)\}},~{}~{}~{}~{}-\infty<y<\infty,~{}0\leq\lambda<\infty.$
The parameter $\lambda$ can be interpreted in terms of the ratio of the hazard rates between $G$ and $F$, as expressed by the following equation:
@@
Notably, $G$ encompasses a broader class of distributions compared to the original distribution $F$; we name it compounded family of distributions. The hazard ratio, denoted as $[1+\lambda\bar{F}(y)]$, indicates that the model follows a proportional hazard-based representation similar to the Cox proportional hazard model Cox (1972). Note that, $\lim_{y\to-\infty}r_{G}(y;\lambda)=\displaystyle\lim_{y\to-\infty}(1+\lambda)r_{F}(y)$ and $\lim_{y\to\infty}r_{G}(y;\lambda)=\displaystyle\lim_{y\to\infty}r_{F}(y).$ The following theorem establishes limits on the hazard rate function and the survival function.
@@
The subsequent theorem provides valuable insights into the concepts of excess risk (the rate of disease in an exposed population divided by the rate of disease in an unexposed population) and relative risk (a ratio of the probability of an event occurring in the exposed group versus the probability of the event occurring in the non-exposed group), which will be of interest in epidemiology and medical science.
@@ III.2 Moments
where $a_{j,k}=\frac{1}{j!}\binom{j}{k}(-1)^{j-k}\lambda^{j}$. Probability weighted moments (PWMs), proposed by Greenwood et al. Greenwood et al. (1979), are statistical measures that represent expectations of certain functions of a random variable when its mean exists. If we express the probability-weighted moments in terms of the density function $f(y)$, then the corresponding equation is given by
@@
Therefore, we can see a closed form for the moments of the proposed family of distribution. The following result shows that the mean residual life function of the random variable $Y$ can be represented by PWMs of the residual random variable $X_{t}=\left(X-t|X>t\right)$.
@@
The mean residual life function $\mu_{G}(t)$ at age $t$ is defined to be the expected remaining life given survival to age $t$; it is a function of interest in actuarial studies, survival analysis, and reliability.
@@ IV Proposed CBurr distribution
Irving W. Burr introduced $12$ distinct cumulative distribution functions, which have been valuable for data analysis Burr (1942). Among these functions, the Burr $XII$ distribution (commonly known as the Burr distribution) has garnered significant attention in the past decade due to its ability to generate a broad range of skewness and kurtosis values Tadikamalla (1980). The Burr distribution is particularly well-suited for analyzing heavy-tailed network data, surpassing distributions such as Exponential, Gamma, and Weibull.
@@ IV.1 Definition and Statistical Properties
In the specific case where the random variables $X_{i}$ follow the Burr distribution with the distribution function given in Eq. (11), the proposed compounded family of distribution in Eq. (1) can be written as:
$\bar{G}(y)=e^{-\lambda}\left[1+\left(\frac{y}{\gamma}\right)^{c}\right]^{-\alpha}e^{\lambda\left(1+\left(\frac{y}{\gamma}\right)^{c}\right)^{-\alpha}}.$
@@
Therefore, the mean residual life function can be expressed as
@@
Remark 2
The CBurr distribution has four parameters and is flexible enough to analyze highly skewed and heavy-tailed datasets. As explained in the next subsection, estimating the parameters is not difficult. It can also be naturally extended for multivariate setup which is a great advantage of CBurr distribution in practical usage.
IV.2 Parameter Estimations: MLE
@@
Upon equating the partial derivatives to zero, a system of equations is obtained. Solving this system simultaneously yields the maximum likelihood estimates (MLE) $\hat{\Theta}=(\hat{\lambda},\hat{\beta}^{T})^{T}$ for $\Theta=(\lambda,\beta^{T})^{T}$. Due to the absence of closed-form solutions for these equations, numerical solutions obtained through iterative algorithms are required to solve them. The CBurr distribution exhibits different shapes due to the changing values of parameters, as depicted in Figs. 2 and 3.
Remark 3
These partial derivatives are set to zero and must be solved simultaneously to obtain the MLEs of the parameters of the CBurr model. As the closed-form solutions are not available, we have used an iterative method to solve these equations numerically. We have used the quasi-Newton Broyden Fletcher Goldfarb Shanno (BFGS) algorithm available in the R package ‘optimx’ for solving the unconstrained nonlinear optimization problem. This helps in finding the MLE estimates of the CBurr model parameters.
@@
Figure 3: 3D Plots corresponding to PDFs of CBurr distribution of varying values of parameters $\gamma$ and $\alpha$. (c and $\lambda$ are constant).
@@ V.1 Datasets
The degree distribution of biological networks is an essential factor in comprehending their structure and function. For example, in PPI networks, node degree distribution indicates the presence of densely packed hub proteins and the majority of proteins having a lower degree. In DDI networks, node degree distributions provide insight into the drug interaction pattern and its pharmacological implications. Additionally, metabolic networks display a scale-less distribution, emphasizing the highly connected metabolites essential for metabolic pathways. Finally, GRN networks use node degrees to comprehend the pattern of gene regulation and to identify important regulatory genes. Analyzing node degree distributions provides researchers with valuable insight into the organization principles and regulatory dynamic of biological networks. The following datasets were used for experiments in this study: Bio-Dmela Singh, Xu, and Berger (2008b): PPI networks were constructed for five species, such as baker’s yeast, fruit fly, nematode worm, mouse, and human. The networks were created by combining data from multiple databases. The PPI data coverage varied among the species, with different numbers of edges observed. Singleton nodes were added based on sequence data to include proteins not present in the PPI network. We collected the following biological network datasets from https://networkrepository.com/bio.php.
@@
Figure 4: A visual representation of the (left) bio-HS-CX data and (right) bio-mouse-Gene to provide insight into its structure, interactions, and functional relationships Rossi and Ahmed (2015).

### v3

@@ III Model Formulation
The proposed Compounded Burr (CBurr) model is motivated by the following intuition that node degree results from a random number of latent interactions, each with heavy-tailed potential. For example, if we consider protein-protein interactions (PPI), the degree of a node (i.e., the number of interactions) may result from a random number of underlying biological processes (e.g., gene expression levels, functional complexes, evolutionary duplications), which the Poisson component captures. The Burr distribution is popularly used in statistics and survival analysis for modeling heavy-tailed data effectively. This makes it suitable for the long right tails seen in biological network degrees. The proposed compounded Burr model uses a Poisson-shifting that adds flexible truncation or tempering effect. By taking the minimum of a Poisson-distributed number of Burr variables, the resulting distribution gains heavier or tempered tails, enabling it to mimic real-world degree distributions more closely than Burr alone. This compounded nature allows the model to capture both global variability (from Poisson) and local heavy-tail scaling (from Burr) — ideal for biological networks where some hubs dominate while others have moderate or sparse connectivity.
The rationale behind proposing a CBurr model, where $N-1\sim\operatorname{Poisson}(\lambda)$ with $\lambda\in(0,\infty)$ and $X_{1},X_{2},\cdots\sim$ Burr - lies in capturing both heavy-tail behavior and variability in the number of competing mechanisms in biological networks. Let $Y=\min\{X_{1},X_{2},\cdots,X_{N}\}$ be the effective or observable node (protein in PPI network example) degree, which gives the minimum of multiple potential contributions. Each $X_{i}$ is the degree a node would have under a specific interaction-generating mechanism (e.g., a biological process, experimental condition, structural domain, or evolutionary module), and each follows a Burr distribution. Here, $N$ is the number of such mechanisms the protein is exposed to, or the number of interaction opportunities or modules a protein is involved in, modeled as a Poisson random variable. For example, a protein might participate in a random number of functional domains, biological pathways, or interaction modules. Each of these modules offers an opportunity for the protein to gain connections, so the total number of such opportunities (interactions) is modeled as a Poisson variable plus 1 (since the model requires at least one event), i.e., $N-1\sim\operatorname{Poisson}(\lambda)$. Since $X_{i}$ represents the potential interaction strength or connectivity level contributed by a single latent mechanism, therefore, Burr distribution is used because it can capture heavy-tailed behavior, representing that some mechanisms may yield many interactions (hubs) while most yield few. So, if a protein is part of $N$ such modules, the degree of the node is modeled as the minimum of these Burr-distributed random variables, mimicking competition or limiting influence between mechanisms. Therefore, this compounded framework helps account for over-dispersion, truncation, and tail behavior observed in real-world biological networks. Now, we present this new compounded family of distributions for the general case.
Let $X_{1},X_{2},\ldots$ be i.i.d. (independent and identically distributed) random variables with CDF $F(x)$ and let $Y=\min\left\{X_{1},\ldots,X_{N}\right\},\quad\text{ where }N-1\sim\operatorname{Poisson}(\lambda)\;\text{with}\;\lambda\in(0,\infty),$ where $\lambda$ controls the expected number of potential interaction trials. So, we have a random number of i.i.d. shocks, but there is always at least one shock with $N\in\{1,2,3,\ldots\}$, and
$\mathbb{P}(N=n)=\frac{e^{-\lambda}\lambda^{n-1}}{(n-1)!},n=1,2,\ldots$
This reflects a realistic biological setting where every node (e.g., a protein) has at least one potential interaction. But additional connections or roles (beyond the first) are random, modeled via a Poisson distribution. Then the survival function for $Y$ becomes:
$\bar{G}(y)$ $=P\left[X_{1}\geq y,X_{2}\geq y,\ldots,X_{n}\geq y\right]:=P(Y\geqslant y)$
@@
$=\displaystyle\sum_{n=1}^{\infty}\left[\bar{F}(y)\right]^{n}\frac{e^{-\lambda}\lambda^{n-1}}{(n-1)!}=\frac{\bar{F}(y)}{\exp\{\lambda F(y)\}},$ (1)
where $\bar{F}(y)$ and $\bar{G}(y)$ represent the survival function of $X$ and $Y$. Mathematically, $\bar{F}=1-F$, $\bar{G}=1-G$, and $F(x)=P[X\leq x]$ is the distribution function of $X$. For non-negative random variables, the survival function is more meaningful with a more convenient form than the better-known cumulative distribution function. $G(y)$ is a valid, smooth, unimodal distribution with a tunable tail via the Poison shift parameter $\lambda$. For modeling biological networks, where data has heavy-tailed behavior as discussed in Sec. II, we choose $\bar{F}$ as the survival function of the Burr random variables. Thus, the survival function in III is not a member of the original Marshall-Olkin (MO) family Marshall and Olkin (1997), but could be seen as a Poisson-shifted Generalized MO distribution.
Remark 1
  1. 1.
    Role of $\lambda$: Suppose $F$ is a Burr random variable, then lower value of $\lambda\;(\sim 0.1-1)$ implies that the model behaves close to Burr; tail is heavy but less suppressed. Very high value of $\lambda\;(>5)$ suggests that the tail becomes more sharply truncated $\rightarrow$ (mimics power-law cutoff). The intermediate value of $\lambda$ introduces subtle curvature in the body of the distribution, modeling non-monotonic or multi-slope behavior.
  2. 2.
    Interpretation of $\lambda$ in biological networks: For practical use, $\lambda$ can be tuned to reflect biological heterogeneity, for e.g., networks with more alternative binding modes or greater stochasticity in forming connections will require a larger value of $\lambda$. In the PPI network example or other biological networks, a higher value of $\lambda$ indicates more variability and more trials before observing the minimal interaction event and a lower value of $\lambda$ suggests fewer competing trials, so the minimum may be less extreme. Overall, it helps in mimicking realistic saturation or resource constraints, when the probability of forming new connections diminishes as the degree grows.
It is worth noting that the ratio ${\bar{F}(y)}/{\bar{G}(y)}$ is non-decreasing for $y\geqslant 0$, shown as
@@
We also investigate the distribution of the maximum, i.e., $Z=\max\{X_{1},X_{2},\cdots,X_{N}\}.$ This will be useful when the focus is on rare, highly connected nodes, and tail behavior dominates the network’s structure. We write the distribution function of $Z$ as follows:
@@
Therefore, $G(\cdot)$ is a generalization of the base distribution $F(\cdot)$ and $G$ satisfies all the properties of a cumulative distribution function when $X$ is a continuous random variable with the cumulative distribution function $F(\cdot)$. However, this study focuses on the distribution function defined in (III) using a Poisson-distributed number of Burr shocks and taking the minimum that is more relevant for biological networks.
A skewing mechanism is a method for introducing additional asymmetry or tail variation into a distribution by applying a probabilistic transformation. This compounding method acts as a skewing mechanism, as illustrated by the following theorem:
@@ III.1 Hazard rates and risk functions
In the context of network data analysis, we introduced a transformation to the distribution $G(y:\lambda)$, which incorporates a new parameter $\lambda\geq 0$ into any existing family of continuous distributions. The parameter $\lambda$ can also be interpreted in terms of the ratio of the hazard rates between $G$ and $F$, as expressed by the following equation:
@@
Notably, $G$ encompasses a broader class of distributions compared to the original distribution $F$; we name it compounded family of distributions. The hazard ratio, denoted as $[1+\lambda\bar{F}(y)]$, indicates that the model follows a proportional hazard-based representation similar to the Cox proportional hazard model Cox (1972). Note that
$\lim_{y\to-\infty}r_{G}(y;\lambda)=\displaystyle\lim_{y\to-\infty}(1+\lambda)r_{F}(y)\;\text{and}\;\displaystyle\lim_{y\to\infty}r_{G}(y;\lambda)=\displaystyle\lim_{y\to\infty}r_{F}(y).$
The hazard rate of the compounded model can be interpreted as follows: (a) If $\lambda>0$, then the hazard is inflated in the early range; (b) If $\lambda<0$ (theoretically impossible but, in practice, happens due to poor initialization), then the hazard is deflated, possibly leading to non-monotonic shapes; (c) When $\lambda=0$, the model reduces to the baseline distribution. In the PPI network example, at lower degrees, the hazard rate is typically low, indicating that proteins with few interactions have a reduced propensity to form new links. As node degree increases, the hazard rate tends to rise, reflecting a phase of rapid interaction accrual commonly associated with hub-like or multifunctional proteins. This non-monotonic hazard behavior supports the relevance of the CBurr formulation in capturing nuanced biological constraints observed in empirical PPI networks and other biological networks.
The following theorem establishes limits on the hazard rate function and the survival function.
@@
The subsequent theorem provides valuable insights into the concepts of excess risk and relative risk, which will be of interest in epidemiology and medical science Panja et al. (2023). Excess risk in biological networks quantifies how much more likely a node (protein) is to gain further interactions beyond a baseline, helping to identify emerging hubs. Relative risk compares the likelihood of connectivity between two nodes, offering insights into functional disparity. Together, they aid in pinpointing critical proteins and understanding interaction dynamics in health and disease contexts.
@@ III.2 Moments
where $a_{j,k}=\frac{1}{j!}\binom{j}{k}(-1)^{j-k}\lambda^{j}$. Probability weighted moments (PWMs), proposed by Greenwood et al. Greenwood et al. (1979), are statistical measures that represent expectations of certain functions of a random variable when its mean exists. If we express the PWMs in terms of the density function $f(y)$, then the corresponding equation is given by
@@
Therefore, we can see a semi-closed-form for the moments of the proposed family of distributions (when the baseline distribution is Burr/Pareto distributions). In biological network modeling, even if closed-form moments are unavailable, empirical moments and simulation-based estimates (bootstrap, MCMC) remain powerful tools to quantify mean connectivity, variability, and skewed risk of interaction across different protein classes. The following result shows that the mean residual life (MRL) function of the random variable $Y$ can be represented by PWMs of the residual random variable $X_{t}=\left(X-t|X>t\right)$.
@@
The MRL function $\mu_{G}(t)$ at age $t$ is defined to be the expected remaining life given survival to age $t$; it is a function of interest in actuarial studies, survival analysis, and reliability. The MRL function plays an important role in understanding network growth dynamics in biological systems like PPI networks. In these networks, nodes represent proteins, and edges represent the interactions between them. The MRL quantifies the expected number of additional interactions a protein may gain, given it already has a certain number of connections.
@@ IV Proposed CBurr distribution
Irving W. Burr introduced $12$ distinct cumulative distribution functions, which have been valuable for data analysis Burr (1942). Among these functions, the Burr $XII$ distribution (commonly known as the Burr distribution) has garnered significant attention in the past decade due to its ability to generate a broad range of skewness and kurtosis values Tadikamalla (1980). The Burr distribution is particularly well-suited for analyzing heavy-tailed network data, surpassing distributions such as Exponential, Gamma, and Weibull Chakraborty, Das, and Chattopadhyay (2022).
@@ IV.1 Definition and Statistical Properties
In the specific case where the random variables $X_{i}$ follow the Burr distribution with the distribution function given in Eq. (11), the proposed compounded family of distribution in Eq. (III) can be written as:
$\bar{G}(y)=e^{-\lambda}\left[1+\left(\frac{y}{\gamma}\right)^{c}\right]^{-\alpha}e^{\lambda\left(1+\left(\frac{y}{\gamma}\right)^{c}\right)^{-\alpha}}.$ (15)
@@
The hazard starts small for small $y$ and grows non-monotonically depending on $\gamma,\alpha,c,\lambda$. When $\lambda$ is large, the exponential term dampens the tail, reducing hazard at high $y$. When $\lambda<0$ (theoretically implausible but in practice sometimes observed while fitting), hazard can spike unnaturally, indicating modeling instability or data underfit. The hazard rate in the CBurr Model captures the instantaneous risk of observing a higher connectivity (degree) in a biological network, given that the current degree is $y$. In the context of PPI networks, this reflects how likely a node (protein) is to acquire additional interactions as its degree increases. A non-monotonic hazard shape—rising initially and then flattening or declining—can indicate that proteins with moderate degrees are more likely to gain new interactions, while high-degree nodes may saturate due to biological constraints like spatial limitations or functional specificity. This aligns with empirical observations in real-world biological network data, where hubs exist but are biologically bounded.
@@
For small integer orders (e.g., first and second-order moment), we derived a semi-closed-form using known identities for Burr distribution; however, numerical approximation will be used in practice.
@@
Therefore, the MRL function can be expressed as
@@
For the CBurr Model, the MRL captures not only the heavy-tailed behavior (from the baseline Burr distribution) but also incorporates a decaying interaction probability through the Poisson-shift mechanism. This allows the model to realistically represent biological constraints, such as limited binding capacities or competitive inhibition, which naturally cap the number of possible future interactions.
IV.2 Interpretation of CBurr model parameters
The CBurr distribution exhibits different shapes due to the changing values of parameters, as depicted in Figs. 2. Fig. 2 shows that CBurr distribution captures non-monotonic, heavy-tailed, and skewed behaviors — all within a four-parameter space (one scale parameter, two shape parameters, and one Poisson shift parameter. In the context of the CBurr model, the emergence of non-monotonic shapes in the degree distribution is closely tied to the choice of parameters. Each parameter governs a specific aspect of the distribution’s behavior: (I) The scale parameter ($\gamma$) determines how stretched the distribution is along the $x$-axis; keeping $\gamma$ small (e.g., $0.5\leq\gamma\leq 2$) ensures a sharp early rise in the distribution; (II) The shape parameter ($c$), when set to moderate or high values ($3\leq c\leq 7$), creates sufficient curvature in the mid-range degrees, allowing the model to reflect an initial increase in frequency followed by a decline; (III) The tail sharpness parameter ($\alpha$), within a range like $2\leq\alpha\leq 6$, ensures the distribution doesn’t decay too fast, enabling a flexible transition between the body and the tail; (IV) The Poisson-shift parameter ($\lambda$) introduces a skewing effect; moderate values ($1\leq\lambda\leq 5$) allow the peak to form naturally without flattening or overpowering the body structure. Excessively high $\lambda$ values can suppress non-monotonicity by overemphasizing the tail decay.
These parameter settings are especially relevant when modeling biological networks, where degree distributions often show a peak followed by a heavy tail, reflecting a few highly connected proteins (hubs) alongside many proteins with moderate interactions. The non-monotonic pattern in such networks arises due to functional modularity, evolutionary pressures, and biophysical constraints.
IV.3 Parameter Estimations: Maximum Likelihood Estimation
@@
Upon equating the partial derivatives to zero, a system of equations is obtained. Solving this system simultaneously yields the maximum likelihood estimates (MLE) $\hat{\Theta}=(\hat{\lambda},\hat{\beta}^{T})^{T}$ for $\Theta=(\lambda,\beta^{T})^{T}$. Due to the absence of closed-form solutions for these equations, numerical solutions obtained through iterative algorithms are required to solve them.
These partial derivatives are set to zero and are solved simultaneously to obtain the MLEs of the parameters of the CBurr model. As the closed-form solutions are not available, we have used an iterative method to solve these equations numerically. We have used the quasi-Newton Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm available in the R package ‘optimx’ for solving the unconstrained nonlinear optimization problem. This helps in finding the MLE estimates of the CBurr model parameters.
@@ V.1 Datasets
The degree distribution of biological networks is an essential factor in comprehending their structure and function. For example, in PPI networks, node degree distribution indicates the presence of densely packed hub proteins and the majority of proteins having a lower degree. In DDI networks, node degree distributions provide insight into the drug interaction pattern and its pharmacological implications. Additionally, metabolic networks display a scale-free distribution, emphasizing the highly connected metabolites essential for metabolic pathways. Finally, GRN networks use node degrees to comprehend the pattern of gene regulation and to identify important regulatory genes. Analyzing node degree distributions provides researchers with valuable insight into the organizational principles and regulatory dynamics of biological networks. The following datasets were used for experiments in this study:
Bio-Dmela Singh, Xu, and Berger (2008b): PPI networks were constructed for five species, such as baker’s yeast, fruit fly, nematode worm, mouse, and human. The networks were created by combining data from multiple databases. The PPI data coverage varied among the species, with different numbers of edges observed. Singleton nodes were added based on sequence data to include proteins not present in the PPI network.
@@
Figure 3: A visual representation of the (left) bio-HS-CX data and (right) bio-mouse-Gene to provide insight into its structure, interactions, and functional relationships Rossi and Ahmed (2015). We have used the Interactive visualization platform, called “GraphVis” [^3] https://networkrepository.com/graphvis.php, (Interactive Visual Graph Mining and Machine Learning) for visualizing the graph structures [^4] https://networkrepository.com/Rossi et al. (2018)